# Avoiding the Price of Adaptivity: Inference in Linear Contextual Bandits via Stability

### v1

@@ 5 Experiments
We evaluate the empirical behavior of the proposed penalized–EXP4 algorithm in a stochastic contextual bandit environment with linear loss structure. In this section we provide the simulation details for the standard OLS estimator. Experimental details for the ridge estimator are presented in Appendix B.
@@ 5.1 Simulation Environment
where, $k\in[K]$ and the entries of the weight vector $u_{a,k}$ are i.i.d. draws from $\mathcal{N}(0,12)$ distribution. In our experiments we set $A=K=5$ and $d_{x}=10$.
@@ 5.2 Algorithmic Configuration
$\mathcal{I}^{\mathrm{APS}}_{T}(a):=\big[a^{\top}\widehat{\beta}_{\mathrm{ridge}}-R_{T}\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a},\;a^{\top}\widehat{\beta}_{\mathrm{ridge}}+R_{T}\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a}\big],$
where $R_{T}$ is from (3) with $S=1$. Wald coverage is measured analogously using the confidence interval $\mathcal{I}^{\mathrm{Wald}}_{T}(a)$ from (4) , with $\widehat{\sigma}$ as the sample standard deviation estimate (lai1982least, Lemma 3). For each $T\in\{500,3000\}$, we report empirical coverage of $\mathcal{I}^{\mathrm{APS}}(a)$ vs $\mathcal{I}^{\mathrm{Wald}}(a)$, and their average width.
@@ 5.3 Simulation Plots
$T\in\{500,3000\},\qquad A=K=5,\qquad d_{x}=10,$
Figure 1: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=500$ runs.
Figure 2: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=3000$ runs.
@@
Figure 4: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=500$ runs.
Figure 5: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=3000$ runs.
@@ 6 Proof of Theorems
In this section we prove Theorems 1 and 2. Recall that $S_{T}=\sum^{T}_{t=1}c(x_{t},a_{t})c(x_{t},a_{t})^{\top}$ where, the random vector $c(x,a)\in^{d}$ satisfies
$\sup_{x,a}\|c(x,a)\|_{2}<\infty.$ (28)
@@
$\mathbb{E}[Y_{t}\mid\mathcal{F}_{t-1}]=\mathbb{E}_{x_{t}}\!\left[\sum_{a\in\mathcal{A}}Q_{t}(a\mid x_{t})\,c(x_{t},a)c(x_{t},a)^{\top}\right]=\sum^{K}_{k=1}w_{t,k}\Sigma_{k}$
@@ 6.1 Proof of Theorem 1
We first prove stability holds for the standard OLS estimator, assuming that $S_{T}$ is invertible, followed by the proof for the modified ridge estimator. Once the stability condition (1) is verified, the asymptotic normality claim (26a) follows directly by invoking Theorem 3 from lai1982least along with an application of Slutsky’s theorem.
@@
From equations (6) and (34), we arrive at the following lower bound:
@@
Part (b) : Proof of stability condition for ridge estimator
$\mathcal{R}_{1}(T):=\dfrac{a^{\top}(\hat{\beta}_{rid}-\beta^{\star})}{\sqrt{a^{\top}R_{T}^{-1}a}}$ (40)
$\tilde{\Sigma}^{-1}_{T}R_{T}\xrightarrow{\mathbb{P}}\mathbf{I}$ (41)
$\mathcal{R}_{1}(T)=\underbrace{\dfrac{a^{\top}(R_{T}^{-1}S_{T}-\mathbf{I})}{\sqrt{a^{\top}R_{T}^{-1}a}}\beta^{\star}}_{\mathcal{T}_{1}(T)}\ \ +\ \ \underbrace{\dfrac{a^{\top}R_{T}^{-1}}{\sqrt{a^{\top}R_{T}^{-1}a}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}}_{\mathcal{T}_{2}(T)}$ (43)
$\mathcal{T}_{1}(T)=-\lambda_{rid}\dfrac{a^{\top}R_{T}^{-1}}{\sqrt{a^{\top}R_{T}^{-1}a}}\beta^{\star}$
@@
$\mathcal{T}_{1}(T)=\frac{-\lambda_{rid}}{\sqrt{T}}\dfrac{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}}{\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}}\beta^{\star}+\frac{-\lambda_{rid}}{\sqrt{T}}\dfrac{a^{\top}[\mathcal{M}_{T}-\mathbf{I}]\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}}{\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}}\beta^{\star}$
Now, for $\mathcal{T}_{2}(T)$ we have a similar decomposition:
@@
$\mathcal{T}_{2}(T)=\underbrace{\dfrac{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}}{\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}}\frac{1}{\sqrt{T}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}}_{I_{1}(T)}\ \ +\ \ \underbrace{\dfrac{a^{\top}[\mathcal{M}_{T}-\mathbf{I}]\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}}{\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}}\frac{1}{\sqrt{T}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}}_{I_{2}(T)}$
Now, from central limit theorem for martingale difference sequence (dvoretzky1972asymptotic), $I_{1}(T)\xrightarrow{d}\mathcal{N}(0,1)$ . This result, combined with the fact that $\mathcal{M}_{T}\xrightarrow{\mathbb{P}}\mathbf{I}$ implies $I_{2}(T)$ converges to 0 in probability. This completes the proof.

### v2

@@ 5 Experiments
We evaluate the empirical behavior of the proposed regularized–EXP4 algorithm in a stochastic contextual bandit environment with linear loss structure. In this section we provide the simulation details for the standard OLS estimator. Experimental details for the ridge estimator are presented in Appendix B.2.
@@ 5.1 Simulation Environment
Extension of the feature map and parameter space.
To ensure that the losses are non-negative, we extend the original feature representation by introducing an intercept term. For each context–action pair $(x_{t},a_{t})$, the original feature vector
$c(x_{t},a_{t})\in\mathbb{R}^{Ad_{x}}$
is augmented as
$\tilde{c}(x_{t},a_{t})$ $=\begin{pmatrix}c(x_{t},a_{t})\\ 1\end{pmatrix}\in\mathbb{R}^{Ad_{x}+1}.$
Correspondingly, the unknown parameter vector is extended to
$\tilde{\beta}^{\star}$ $=\begin{pmatrix}\beta^{\star}\\ 2\end{pmatrix}\in\mathbb{R}^{Ad_{x}+1},$
Under this augmented representation, the loss model becomes
$\ell_{t}$ $=\langle\tilde{c}(x_{t},a_{t}),\tilde{\beta}^{\star}\rangle+\varepsilon_{t},$
which is algebraically equivalent to the original linear model but explicitly accounts for a constant offset in the losses. In particular, as $|\varepsilon_{t}|\leq 0.1,$ and $|c(x_{t},a_{t})|\leq 1$ adding shift of constant $2$ to the observed loss ensures that the resultant losses are non-negative.
@@
where, $k\in[K]$ and the entries of the weight vector $u_{a,k}$ are i.i.d. draws from $\mathcal{N}(0,0.04)$ distribution. In our experiments we set $A=K=5$ and $d_{x}=10$.
@@ 5.2 Algorithmic Configuration
$\mathcal{I}^{\mathrm{APS}}_{T}(a):=\big[a^{\top}\widehat{\beta}_{\mathrm{ridge}}-\xi_{T}\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a},\;a^{\top}\widehat{\beta}_{\mathrm{ridge}}+\xi_{T}\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a}\big],$
where $\xi_{T}$ is
$\xi_{T}:=\sqrt{2\!\left(\frac{1}{2}\log\frac{\det(V_{t})}{\det(\lambda I)}\;+\;\log\frac{1}{\alpha}\right)}\;+\;\sqrt{\lambda}\,\|\beta^{\star}\|_{2}$ (28)
We note that the confidence intervals $\mathcal{I}^{\mathrm{APS}}_{T}(a)$ defined above are sharper than the anytime valid confidence interval defined in equation (2). Wald coverage is measured analogously using the confidence interval $\mathcal{I}^{\mathrm{Wald}}_{T}(a)$ from (4) , with $\widehat{\sigma}$ as the sample standard deviation estimate (lai1982least, Lemma 3). For each $T\in\{500,3000\}$, we report empirical coverage of $\mathcal{I}^{\mathrm{APS}}(a)$ vs $\mathcal{I}^{\mathrm{Wald}}(a)$, and their average width.
@@ 5.3 Simulation Plots
$T\in\{500,3000\},\qquad A=K=5,\qquad d_{x}=30,$
Figure 1: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. The average CI widths of $\mathcal{I}^{\mathrm{Wald}}$ and $\mathcal{I}^{\mathrm{APS}}$ across all values of $\alpha$ are $0.08$ and $2.61$ respectively. Simulations are based on $T=500$ runs.
Figure 2: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. The average CI widths of $\mathcal{I}^{\mathrm{Wald}}$ and $\mathcal{I}^{\mathrm{APS}}$ across all values of $\alpha$ are $0.03$ and $1.15$ respectively. Simulations are based on $T=3000$ runs.
@@
Figure 4: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. The average CI widths of $\mathcal{I}^{\mathrm{Wald}}$ and $\mathcal{I}^{\mathrm{APS}}$ across all values of $\alpha$ are $0.12$ and $4.55$ respectively. Simulations are based on $T=500$ runs.
Figure 5: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. The average CI widths of $\mathcal{I}^{\mathrm{Wald}}$ and $\mathcal{I}^{\mathrm{APS}}$ across all values of $\alpha$ are $0.06$ and $1.88$ respectively. Simulations are based on $T=3000$ runs.
@@ 6 Proof of Theorems
In this section we prove Theorems 1 and 2. Recall from Assumption (A2) that $S_{T}=\sum^{T}_{t=1}c(x_{t},a_{t})c(x_{t},a_{t})^{\top}$ where, the random vector $c(x,a)\in^{d}$ satisfies
$\sup_{x,a}\|c(x,a)\|_{2}\leq 1.$ (29)
@@
$\mathbb{E}[Y_{t}\mid\mathcal{F}_{t-1}]=\mathbb{E}_{x_{t}}\!\left[\sum_{a\in\mathcal{A}}Q_{t}(a\mid x_{t})\,c(x_{t},a)c(x_{t},a)^{\top}\right]=\sum^{K}_{k=1}w_{t,k}\Sigma_{k}$ (31)
@@ 6.1 Proof of Theorem 1
We first prove stability holds for the standard OLS estimator, assuming that $S_{T}$ is invertible. Once the stability condition (1) is verified, the asymptotic normality claim follows directly by invoking Theorem 3 from lai1982least along with an application of Slutsky’s theorem.
@@
The above inequality is justified as on $\Delta_{\varepsilon}$, the Bregman divergence induced by the (shifted) negative entropy satisfies $D_{R}(p,q)=\mathrm{KL}(p\|q)$ (the additive and linear terms cancel on the simplex). From equations (6) and (36), we arrive at the following lower bound:
@@
Part (b): Proof of the quantitative CLT
The standard approach to prove such Berry Esseen bound is to uncover and utilize a martingale difference structure in our statistic of interest (hall2014martingale). For notational simplicity, we assume that the sample covariance matrix $S_{T}$ is invertible. If $S_{T}$ is not invertible, a modified ridge estimator using $R_{T}:=S_{T}+\lambda_{rid}I$ exhibits equivalent asymptotic behavior. We highlight the details in Appendix B.
Fix a vector $a\in^{d}\setminus\{0\}$ and define
@@
$\mathcal{S}_{1}(T):=\frac{a^{\top}\left(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star}\right)}{\sqrt{a^{\top}S_{T}^{-1}a}}=\frac{a^{\top}S_{T}^{-1}}{\sqrt{a^{\top}S_{T}^{-1}a}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}.$ (43)
where $z_{t}:=c(x_{t},a_{t})$ .We note that since $S_{T}$ is a random matrix measurable with respect to $\mathcal{F}_{T}$, and $\mathcal{S}_{1}(T)$ is not a sum of martingale difference sequence.
Define $\tilde{\Sigma}_{T}:=\mathbb{E}[S_{T}]$ and consider the alternate statistic
$\mathcal{S}_{2}(T):=\frac{a^{\top}(\tilde{\Sigma}_{T})^{-1}}{\sqrt{a^{\top}(\tilde{\Sigma}_{T})^{-1}a}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}$ (44)
Observe that assumption (A4) ensures that $\tilde{\Sigma}_{T}$ is invertible. Let $b=(\tilde{\Sigma}_{T})^{-1/2}\ a$. Algebraic manipulation yields that,
$\mathcal{S}_{2}(T)=\frac{b^{\top}(\tilde{\Sigma}_{T})^{-1/2}}{\|b\|_{2}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}=\frac{1}{\|b\|_{2}}\sum^{T}_{t=1}h_{t}\varepsilon_{t}$
where $h_{t}:=b^{\top}(\tilde{\Sigma}_{T})^{-1/2}z_{t}$. We observe that
$\mathbb{E}[h_{t}\varepsilon_{t}\mid\mathcal{F}_{t-1}]$ $=\mathbb{E}[\mathbb{E}[h_{t}\varepsilon_{t}\mid\mathcal{F}_{t-1},x_{t},a_{t}]\mid\mathcal{F}_{t-1}]$
$=\mathbb{E}[h_{t}\mathbb{E}[\varepsilon_{t}\mid\mathcal{F}_{t-1},x_{t},a_{t}]\mid\mathcal{F}_{t-1}]$
$=0$
Therefore, $(h_{t}\varepsilon_{t},\mathcal{F}_{t})_{t\geq 1}$ is a martingale difference sequence and one may analyze $\mathcal{S}_{2}(T)$ using standard martingale CLT (mourrat2013rate). Letting $\mathcal{M}_{T}(p):=k_{T}S_{T}^{-1}\tilde{\Sigma}_{T}$ where $k_{T}=\sqrt{a^{\top}\left(\tilde{\Sigma}_{T}\right)^{-1}a/a^{\top}S_{T}^{-1}a}$ we may rewrite our statistic of interest $\mathcal{S}_{1}(T)$ in terms of $\mathcal{S}_{2}(T)$:
@@
$\mathcal{S}_{1}(T)=\mathcal{S}_{2}(T)\ +\frac{a^{\top}\left(\mathcal{M}_{T}(p)-\mathbf{I}\right)}{\|b\|_{2}}\sum^{T}_{t=1}\left(\tilde{\Sigma}_{T}\right)^{-1}z_{t}\varepsilon_{t}=\mathcal{S}_{2}(T)+\mathcal{V}_{T}$ (45)
where $\mathcal{V}_{T}:=\frac{a^{\top}\left(\mathcal{M}_{T}(p)-\mathbf{I}\right)}{\|b\|_{2}}\sum^{T}_{t=1}\left(\tilde{\Sigma}_{T}\right)^{-1}z_{t}\varepsilon_{t}$. In Lemma 3 we argue that the behavior of $\mathcal{S}_{1}(T)$ and $\mathcal{S}_{2}(T)$ are equivalent for large $T$. Formally, we have
Lemma 3.
Let $\mathcal{N}_{T}:=\left[\frac{1}{T}\tilde{\Sigma}_{T}\right]^{-1}\times\left[\frac{M_{T}}{T}+\frac{B_{T}}{T}-\frac{1}{T}\tilde{\Sigma}_{T}\right]$ with $\tilde{\Sigma}_{T}=\mathbb{E}[S_{T}]$, and $Z$ be a copy of the standard normal random variable. Suppose that
$\quad\operatorname{d}_{\operatorname{K}}\left(\mathcal{S}_{2}(T),Z\right)\leq\beta_{T}$ (46)
where $\beta_{T}$ is some positive real sequence. Then,
$\operatorname{d}_{\operatorname{K}}\left(\mathcal{S}_{1}(T),Z\right)\lesssim\beta_{T}+2\cdot\Psi(\gamma_{T})^{1/2}\ +\ d\cdot\operatorname{exp}\left\{-\frac{(\lambda^{\star}_{L})^{2}}{32+8\lambda^{\star}_{L}/3}\cdot T\right\}$ (47)
where $\Psi(\gamma_{T})$ as defined in equation (25b).
It now remains to analyze the term $\mathcal{S}_{2}(T)$, and we do so by using a Berry-Esseen bounds for martingale CLTs (mourrat2013rate). Lemma 4 stated below characterizes $\beta_{T}$.
Lemma 4.
Let $a\in^{d}$ be any arbitrary real vector, $\tilde{\Sigma}_{T}=\mathbb{E}[\sum^{T}_{t=1}z_{t}z^{\top}_{t}]$. Then
$\operatorname{d}_{\operatorname{K}}\left(\mathcal{S}_{2}(T),Z\right)\leq\ C\left[\Psi(\gamma_{T})^{1/3}\ +\ \frac{1}{T^{1/3}}\right]$ (48)
where $\Psi(\gamma_{T})$ as is defined in equation (25b) and $C$ is a constant independent of $T$.
Hence by combining equations (47) and (48) together we obtain:
$\operatorname{d}_{\operatorname{K}}\left(\mathcal{S}_{1}(T),Z\right)\lesssim\Psi(\gamma_{T})^{1/3}\ +\ \frac{1}{T^{1/3}}\ +\ 2\cdot\Psi(\gamma_{T})^{1/2}\ +\ d\cdot\operatorname{exp}\left\{-\frac{(\lambda^{\star}_{L})^{2}}{32+8\lambda^{\star}_{L}/3}\cdot T\right\}$
As $\Psi(\gamma_{T})^{1/2}$ is strictly dominated by $\Psi(\gamma_{T})^{1/3}$, we may ignore this term. This completes the proof of Theorem 1. We prove Lemma 4 in Appendix A, and prove our key Lemma 3 next.
Proof of Lemma 3:
Recall the following decomposition.
$\mathcal{S}_{1}(T)=\mathcal{S}_{2}(T)\ +\frac{a^{\top}\left(\mathcal{M}_{T}(p)-\mathbf{I}\right)}{\|b\|_{2}}\sum^{T}_{t=1}\left(\tilde{\Sigma}_{T}\right)^{-1}z_{t}\varepsilon_{t}=\mathcal{S}_{2}(T)+\mathcal{V}_{T}$ (49)
where $\mathcal{V}_{T}:=\frac{a^{\top}\left(\mathcal{M}_{T}(p)-\mathbf{I}\right)}{\|b\|_{2}}\sum^{T}_{t=1}\left(\tilde{\Sigma}_{T}\right)^{-1}z_{t}\varepsilon_{t}$ and $\mathcal{M}_{T}(p):=k_{T}S_{T}^{-1}\tilde{\Sigma}_{T}$ such that
$k_{T}=\sqrt{\frac{a^{\top}\left(\tilde{\Sigma}_{T}\right)^{-1}a}{a^{\top}S_{T}^{-1}a}}$
Now, fix a positive sequence $\delta_{1,T}$ whose explicit choice will be made later in the proof. Define event $\mathcal{E}_{1}(T)$ such that
$\mathcal{E}_{1}(T):=\left\{|\mathcal{V}_{T}|\leq\delta_{1,T}\right\}$ (50)
On event $\mathcal{E}_{1}(T)$, we have
$\mathcal{S}_{2}(T)-\delta_{1,T}\leq\mathcal{S}_{1}(T)\leq\mathcal{S}_{2}(T)+\delta_{1,T}$ (51)
Furthermore, suppose that for every sequence $\delta_{1,T}$, there exists another sequence $\delta_{2,T}$ such that
$\mathbb{P}(\mathcal{E}_{1}(T)^{c})\leq\delta_{2,T}$ (52)
Now we are ready to prove our claim (47), which we prove in two steps. We use equation (51) to derive lower and upper bounds on $\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)$, which are free of $x$.
$\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)$ $\geq\mathbb{P}(\mathcal{S}_{1}(T)\leq x,\mathcal{E}_{1}(T))-\Phi(x)$
$\geq\mathbb{P}(\mathcal{S}_{2}(T)+\delta_{1,T}\leq x,\mathcal{E}_{1}(T))-\Phi(x)$
$\geq\mathbb{P}(\mathcal{S}_{2}(T)\leq x-\delta_{1,T})-\Phi(x)-\mathbb{P}(\mathcal{E}_{1}(T)^{c})$
From the last inequality we have:
$\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)$
$\geq\bigg[\underbrace{\mathbb{P}(\mathcal{S}_{2}(T)\leq x-\delta_{1,T})-\Phi(x-\delta_{1,T})}_{\mathcal{I}_{1}}\bigg]+\bigg[\underbrace{\Phi(x-\delta_{1,T})-\Phi(x)}_{\mathcal{I}_{2}}\bigg]-\underbrace{\mathbb{P}(\mathcal{E}_{1}(T)^{c})}_{\mathcal{I}_{3}}$
Now, from our assumption (46) we have $\mathcal{I}_{1}\geq-\beta_{T}$. By applying first order Taylor expansion along with the fact that $\sup_{x}e^{-x^{2}}<1$ leads us to $\mathcal{I}_{2}\gtrsim-\delta_{1,T}$. Finally, equation (52) implies that $\mathcal{I}_{3}\leq\delta_{2,T}$. Hence, for all $x\in$,
$\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)\gtrsim-\bigg[\beta_{T}+\delta_{2,T}+\delta_{1,T}\bigg]$ (53)
Now, for the upper bound, observe that
$\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)$ $=\mathbb{P}(\mathcal{S}_{1}(T)\leq x,\mathcal{E}_{1}(T))+\mathbb{P}(\mathcal{S}_{1}(T)\leq x,\mathcal{E}_{1}(T)^{c})-\Phi(x)$
$\leq\mathbb{P}(\mathcal{S}_{2}(T)-\delta_{1,T}\leq x,\mathcal{E}_{1}(T))+\mathbb{P}(\mathcal{E}_{1}(T)^{c})-\Phi(x)$
$\leq\mathbb{P}(\mathcal{S}_{2}(T)\leq x+\delta_{1,T})-\Phi(x)+\mathbb{P}(\mathcal{E}_{1}(T)^{c})$
The last inequality leads to,
$\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)$
$\leq\bigg[\underbrace{\mathbb{P}(\mathcal{S}_{2}(T)\leq x+\delta_{1,T})-\Phi(x+\delta_{1,T})}_{\mathcal{I}_{4}}\bigg]+\bigg[\underbrace{\Phi(x+\delta_{1,T})-\Phi(x)}_{\mathcal{I}_{5}}\bigg]+\underbrace{\mathbb{P}(\mathcal{E}_{1}(T)^{c})}_{\mathcal{I}_{3}}$
Analogous calculations yield :
$\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)\lesssim\beta_{T}+\delta_{1,T}+\delta_{2,T}$ (54)
Therefore,
$\sup_{x\in}|\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)|\lesssim\beta_{T}+\delta_{1,T}+\delta_{2,T}$
Define $\lambda^{\star}_{L}$ and $\lambda^{\star}_{U}$ such that
$\min_{k\in[K]}\lambda_{\min}(\Sigma_{k})=\lambda^{\star}_{L}\quad\text{and}\quad\max_{k\in[K]}\lambda_{\max}(\Sigma_{k})=\lambda^{\star}_{U}.$ (55)
Now, the lemma stated below, characterizes $\delta_{2,T}$ in terms of $\delta_{1,T}$, $\lambda^{\star}_{L}$ and $\lambda^{\star}_{U}$.
Lemma 5.
Under the setup of Theorem 1 we have:
$\mathbb{P}(|\mathcal{V}_{T}|>\delta_{1,T})\lesssim\frac{\Psi(\gamma_{T})}{\delta_{1,T}}\ +\ d\cdot\operatorname{exp}\left\{-\frac{(\lambda^{\star}_{L})^{2}}{32+8\lambda^{\star}_{L}/3}\cdot T\right\}$ (56)
We prove this lemma in Appendix A. We choose $\delta_{1,T}$ such that the above upper bound gets minimised. By applying the AM-GM inequality, we obtain the desired bound by choosing $\delta_{T}=\sqrt{\Psi(\gamma_{T})}$ and $\beta_{T}=\Psi(\gamma_{T})^{1/3}$:
$\sup_{x\in}|\mathbb{P}(\mathcal{S}_{1}(T)\leq x)-\Phi(x)|\lesssim\beta_{T}+2\ \Psi(\gamma_{T})^{1/2}\ +\ d\cdot\operatorname{exp}\left\{-\frac{(\lambda^{\star}_{L})^{2}}{32+8\lambda^{\star}_{L}/3}\cdot T\right\}$
Note that $\Psi(\gamma_{T})^{1/2}$ is strictly dominated by $\Psi(\gamma_{T})^{1/3}$, and hence we ignore this term in the upper bound. Therefore, our proof is complete.